# LaDTalk: Latent Denoising for Synthesizing Talking Head Videos with High Frequency Details

### v2

@@ 3 Latent Denoising Modeling
Figure 3: (a) displays the overview of vanilla Wav2Lip[1]. (b) displays the latent denosing process of SOVQAE and the mechanism of regularization loss. Please zoom in for better visualization.
4 LaDTalk
In this section, we introduce the detailed architecture of our LaDTalk. As illustrated in Fig. 3, LaDTalk contains two parts: a vanilla Wav2Lip model synthesis lip-synchronized audio-driven LQ talking faces; an optimized VQAE temporal-consistently lifts LQ faces to HQ faces.
4.1 Vanilla Wav2Lip
As shown in Fig. 3(a), Wav2Lip[1] generates accurate lip-sync by learning from a well-trained SyncNet, where aligns the generated low-half faces and audio in semantic space.
SyncNet processes five-frame face sequences (focusing on the lower half) and the mel-spectrogram of a 0.2-second audio segment, given that the training videos are recorded at 25 FPS. The architecture of SyncNet comprises one audio encoder and one face encoder, both of which are stacks of 2D convolutional neural networks. SyncNet is trained to discern the synchrony between the input audio and video, thereby aiding the lip generator training in a contrastive learning framework.
To accomplish this, training video segments are stochastically sampled to be either in sync or out of sync with audio segments. The training of SyncNet employs a cosine similarity metric coupled with a binary cross-entropy loss, denoted as $\mathcal{L}_{\text{syncnet}}$. Specifically, let $\mathbf{v}\in\mathbb{R}^{D}$ represent the face embedding and $\mathbf{a}\in\mathbb{R}^{D}$ represent the audio segment embedding. The cosine similarity is initially calculated as
${\rm sim}(\mathbf{v},\mathbf{a})=\frac{(\mathbf{v}\cdot\mathbf{a})}{\|\mathbf{v}\|_{2}\|\mathbf{a}\|_{2}},$ (7)
where ()($\cdot$)( ⋅ ) signifies the dot product of vectors. Subsequently, the binary cross-entropy loss is defined as
$\mathcal{L}_{\text{syncnet}}=-y\cdot\log({\rm sim}(\mathbf{v},\mathbf{a}))+(1-y)\cdot\log(1-{\rm sim}(\mathbf{v},\mathbf{a})),$ (8)
where ${\rm sim}(\mathbf{v},\mathbf{a})$ denotes the likelihood that the input audio-video pair is synchronized. A trained SyncNet is then utilized to impose a penalty on the lip generator for any inaccurate generation during the training phase.
The Lip Generator, depicted in Fig. 3(a), comprises three main components: an audio encoder $E_{a}$, a face encoder $E_{f}$, and a face decoder $D_{f}$. The audio encoder $E_{a}$ converts mel-spectrograms of audio into audio features $f_{\text{audio}}\in\mathbb{R}^{D}$. The face encoder $E_{f}$ encodes five reference face frames and five pose face frames into face features $f_{\text{face}}\in\mathbb{R}^{D}$. The reference faces encapsulate essential identity information, such as teeth, lips, and other textures, while the pose faces define the head poses. It is important to note that the lower halves of the pose faces are masked to prevent misleading information regarding the lip region. The unmasked pose faces are the ground truth faces that align with the input audio segment. Additionally, the intermediate features from $E_{f}$ are integrated into the face decoder $D_{f}$ through residual connections [38]. The lowest-level feature of $D_{f}$ is the concatenation of $f_{\text{face}}$ and $f_{\text{audio}}$.
The generator is trained to minimize the L1 reconstruction loss between the generated frames $\mathbf{F}_{\text{gen}}$ and the ground-truth frames $\mathbf{F}_{\text{GT}}$, given by:
$\mathcal{L}_{\text{recon}}=\|\mathbf{F}_{\text{gen}}-\mathbf{F}_{\text{GT}}\|_{1}.$ (9)
Furthermore, to leverage the pretrained SyncNet, five continuous lower half faces and the corresponding 0.2-second audio segment are fed into the pretrained SyncNet. This process minimizes the following expert sync-loss:
$\mathcal{L}_{\text{sync}}=\mathcal{L}_{\text{syncnet}|y=1},$ (10)
where $\mathcal{L}_{\text{sync}}$ ensures that the input audio-video pair is synchronized. This contrastive learning strategy has proven to be effective. According to the statistics from recent papers [2, 14, 3, 39], Wav2Lip continues to achieve the best lip-synchronization performance.
4.2 Space Optimized VQAE
The primary limitations of Wav2Lip are its low resolution ($96\times 96$) and the associated blurring effects. To address these issues, the SOVQAE is designed to consistently upscale LQ faces to HQ faces. Consequently, the SOVQAE should act as a lossless compression technique for talking faces, equipped with a robust noise tolerance capability.
To overcome the resolution limitation, we employ face detection using S3FD [45] to extract facial regions from a talking head video of a specific identity. These extracted regions are subsequently resized and used to train the VQAE. This method allows the codebook to concentrate on capturing the high-frequency facial components. In inference, for given LQ faces, we upscale them to $256\times 256$ before subjecting them to the trained VQAE for latent denoising.
@@ 4.2 Space Optimized VQAE
According to Theorem 3.3, perfect latent denoising is achievable when $\|\mathcal{N}\|_{F}<\frac{d_{C}-2\gamma}{2L_{\epsilon}}$. Note that $\gamma$ is much close to zero, because $\gamma$ is optimized via VQ loss (see Eq. 12) in training. Hence, it may seem optimal to maximize $d_{C}$ while minimizing $L_{\epsilon}$. However, this is theoretically infeasible because $L_{\epsilon}$ is positively correlated with the operation norm of convolutional kernels, implying that minimizing $L_{\epsilon}$ equates to L2/L1 parameter regularization[40]. Such regularization would globally compress the latent space distribution, consequently compressing the distribution of anchors in the codebook $\mathbf{C}$ and reducing $d_{C}$ and degrade the noise robustness.
@@
The collect optimization is that increase $d_{C}$ while maintaining $L_{\epsilon}$. In experiments, we find a practical solution is to locally extend the minimum $d_{C}$ with a given lower bound. Thus, the codebook regularization loss we propose is:
@@
where $\theta$ represents the given upper bound. This regularization strategy effectively improves local noise robustness while maintaining global robustness, which will be validated via ablation studies. In addition to this regularization loss, the complete loss function of SOVQAE comprises L2 reconstruction loss, VQ loss[21], perceptual loss, and GAN loss with a patch-based discriminator[58]. The VQ loss is following:
@@
where $x$ is the input image, $\hat{x}=\delta(g_{\textbf{C}}(\epsilon(x)))$ is the reconstructed image, $\text{sg}(\cdot)$ denotes the stop-gradient operation following[21]. The later two components of  Eq. 12 are utilized to optimized the $\gamma$.
@@ 5 Experiments
Dataset. For comparative analysis, we employ the same set of meticulously edited video sequences utilized in previous works [3, 39, 14], encompassing both English and French dialogues. The average duration of these videos is approximately 8,843 frames. We refer to this collection as the Usual Dataset to differentiate it from our proprietary HF videos. These videos are characterized by their rich textures and high resolution. The average length of these videos is around 11,000 frames. More details of them can be found in supplemental materials. All videos are centered on individual characters and captured at a frame rate of 25 FPS. Consistent with prior research, we adopt a train-to-test split ratio of $10:1$. We empirically set $\theta=1$.
@@
Comparison Baselines. We conduct a comparative analysis of our method with a total of three one-to-all approaches, which include Wav2Lip [1], DINet [12], IP-LAP [51], and three one-to-one NeRF-based methods, namely SyncTalk [2],GeneFace [14] and GeneFace++ [52]. Furthermore, to underscore the significance of our identity-specific face restoration concept, we establish two post-processing baselines using Codeformer [20] and GFPGAN [36].
@@
Evaluation Metrics. To assess the quality of the generated images, we employ the Peak Signal-to-Noise Ratio (PSNR) and the Frechet Inception Distance (FID) [46] as metrics. For evaluating lip synchronization, we utilize the Lip Sync Error Confidence (LSE-C) and Lip Sync Error Distance (LSE-D) scores. Furthermore, to assess the generalizability of our methods, we conduct additional tests using a set of out-of-domain (OOD) audio. This OOD audio comprises 10 diverse voice samples, spanning various languages and genders.
Sliding Evaluation. In our experiments, we observe that the generated video lengths from various methods [1, 12, 14, 2, 39] are not only disparate but also consistently shorter than those of the ground truth videos. Directly assessing image quality in conjunction with video duration could potentially diminish the perceived performance of these methods. To address this issue, we introduce a Sliding Evaluation technique. As illustrated in  Fig. 4, this approach treats the generated video sequence as a dynamic sliding window, calculating the average image quality across all frames captured within the window. The window progressively moves along the timeline, and the maximum value obtained from these iterations is adopted as the final evaluation outcome. We apply this Sliding Evaluation method to all PSNR and FID results to evaluate different methods, thus ensuring a more holistic and equitable assessment.
Figure 4: Illustration of our Sliding Evaluation method.
@@ 5.1 Quantitative Comparison
For a more exhaustive evaluation, our quantitative assessment is divided into two segments: 1) Video quality and lip-synchronization comparison, which evaluates image quality frame-by-frame and lip-synchronization performance under original audio-driven conditions, following prior researches [2, 14]; 2) Out-of-Distribution (OOD) audio-driven comparison, where we contrast lip-synchronization with SOTA methods to underscore the superior cross-audio generalizability of our approach, a critical application scenario. The former assesses identity preservation, while the latter evaluates audio generalization capabilities.
Table 1: The quantitative comparison of different methods on Usual Dataset and HF videos. * indicates post-process baselines. Please zoom in for better visualization.
Method Usual Dataset HF videos
@@
Wav2Lip[1] 34.0979 10.8503 9.077 5.8769 34.6593 17.5281 8.8496 6.8712
DINet[12] 33.9108 9.8385 7.1951 7.4343 34.4849 10.5749 8.3541 7.2340
IP-LAP[51] 34.7263 10.1703 5.1736 8.8160 35.1595 9.9438 8.2604 7.1838
GeneFace[14] 24.8165 21.7084 5.195 - - - - -
GeneFace++[52] 31.1164 20.5506 6.8916 7.5014 33.9692 22.3928 4.6112 11.0401
SyncTalk[2] 36.0574 6.4855 7.054 7.282 39.4050 6.5370 7.816 7.715
CodeFormer*[20] 33.2441 26.3567 9.1739 5.7720 33.7958 16.7215 7.9409 7.2582
GFPGAN*[36] 33.9803 16.7267 9.0417 5.8144 33.7780 17.2715 8.7940 6.5193
Ours 38.9738 3.4505 8.1274 6.44 39.5454 4.3065 8.1392 6.8438
@@
Comparison with End-to-End Methods. As detailed in  Tab. 1, on the Usual Dataset, our method secures the highest scores for both PSNR (38.9738) and FID (3.4505), signifying that we achieve state-of-the-art image quality for generated videos. This advancement is credited to the SOVQAE, which retains high-frequency texture details in the codebook and consistently recovers facial textures from low-quality faces. In terms of lip-synchronization, we outperform all methods (excluding our base model), a testament to Wav2Lip’s robust audio-lip alignment capability. Although our method moderately reduces the lip-synchronization performance of the base model, the substantial enhancement in visual quality justifies this trade-off as valuable and insightful. Similar outcomes are observed in the HF videos.
Comparison with Post-Process Methods. For a comprehensive analysis, we also juxtapose our method with recent blind face restoration SOTA techniques. As illustrated in  Tab. 1, our method surpasses two post-process baselines in video quality. This advantage stems from the fact that Codeformer and GFPGAN are tailored for blind face restoration, and their face prior learning from out-of-distribution multi-face datasets results in information leakage, leading to identity errors. Moreover, their image-by-image training strategy is not optimal for video scenarios, causing frame-wise flickering, which significantly detracts from the final video fidelity. Although they maintain the lip-synchronization performance of the base model, their subpar visual fidelity limits their applicability to talking head scenarios. Furthermore, for fair comparison, we also finetune GFPGAN model in ID-specific videos. As shown in Fig. 5, while fine-tuning improves its performance in visual quality, its synchronization decreases. On the other hand, our method has least GPU need (8.5 GPU@4090 hours), best visual quality and comparable synchronization. These findings strongly affirm the efficacy and efficiency of our post-process approach for talking head tasks.
Method DINet[12] IP-LAP[51] GeneFace++[52] SyncTalk[2] Ours
@@
1 4.7768 8.9890 4.4024 9.3161 4.8974 8.9521 4.2177 9.28129 6.2354 8.0987
2 4.3692 9.0287 4.1531 8.6423 4.6195 8.7908 4.6328 8.0043 5.9480 7.6968
3 5.6343 8.6510 4.8950 9.1588 5.1553 8.8068 6.3866 8.0077 7.4764 7.5445
4 4.2127 8.5808 4.8710 8.6493 3.9984 8.2326 5.7213 7.8372 6.7817 7.5449
5 6.1828 8.5279 3.8994 9.4514 5.9873 8.5823 4.4513 8.8178 6.0246 7.9151
6 6.2382 7.9339 5.1737 8.5354 5.9934 7.7590 6.0401 7.1978 7.1988 7.3960
7 4.6975 8.7730 4.6781 9.2968 5.0774 8.3832 5.4807 8.2488 7.2357 7.1888
8 5.5323 8.3505 4.5502 8.9811 5.9367 7.8090 5.4721 7.9490 6.8894 7.2750
9 5.0672 8.8984 4.7241 9.2790 5.5664 8.6923 5.7529 7.7905 6.9524 7.5649
10 5.5989 8.1889 2.4123 9.2861 6.2121 7.9188 2.9075 7.3401 3.7788 8.4436
Best.
Comparison on OOD Audio Driving. In the industry, voice generalization is crucial for talking head models. While methods like SyncTalk, GeneFace, and GeneFace++ have made progress, their capabilities still fall short of application requirements. To evaluate our method’s voice generalization, we compared it with recent SOTA models on the OOD audio-driving task. As shown in Table 2, Our method leads in both LSE-C and LSE-D metrics, scoring 6.4521 in LSE-C and 7.6668 in LSE-D, demonstrating strong generalization across diverse voices. It significantly outperforms other methods such as SyncTalk, GeneFace++, IP-LAP, and DINet.
Figure 6: In this qualitative analysis, we compare facial synthesis across various methodologies. The comparative visualization presents a curated selection of discrete keyframes and corresponding sub-frames of lower-half faces, each illustrating the performance of different methods on distinct identities. For an optimized visual assessment, we encourage the reader to zoom in on the figures for enhanced clarity. Please zoom in for better visualization.
Figure 7: Comparison on HF videos. Except for the collapse and lip error (see red boxes), NeRF-based methods also has limitation in stability of face scale. See supplemental materials for more details.
These observations suggest that our method effectively inherits the SOTA audio generalization capability of Wav2Lip and validate the feasibility of the Wav2Lip-based post-process approach. Note, we also conduct experiment with different choices of base models in supplemental materials to demonstrate the plug-and-play performance of LaDTalk framework.
5.2 Qualitative Comparison
To provide a more intuitive assessment of image quality, we present a comparative analysis of our method alongside alternative approaches in  Fig. 6. Specifically, we contrast our method with both end-to-end methodologies, including Wav2Lip [1], SyncTalk [2], IP-LAP [51], and DINet [12], as well as two post-processing techniques, GFPGAN [36] and CodeFormer [20].
In Fig. 6, our method outperforms SyncTalk, a state-of-the-art NeRF-based approach, by better capturing high-frequency facial details. SyncTalk’s struggle to maintain these details, particularly in dynamic scenes, results in blurred mouth and teeth regions. This is attributed to NeRF’s limited expressivity for high-frequency details. In contrast, our method, utilizing SOVQAE, effectively restores individual-specific textures, enhancing fidelity. Compared to Wav2Lip, our method not only preserves lip shape but also achieves more realistic facial textures, indicating superior accuracy. The 7th and 8th columns of the figure, where our method’s identity-specific restoration approach is superior to the two post-processing baselines, which alter the subject’s identity due to information leakage from a large OOD faces dataset [47]. This underscores the necessity of our identity-specific restoration approach. Besides, in more challenging HF videos, our method has more stable performance. In contrast, NeRF-based methods occur geometry collapse (see the 3-th cloumn in Fig. 7) and lip error, when scale their resolution to higher.
In summary, the qualitative comparison underscores that our method surpasses all compared methods in terms of facial texture and lip-synchronization, thereby affirming the efficacy of our post-processing strategy. Note that partly qualitative comparison of HF videos can be found in Appendix C.

### v3

@@ 4 Space-Optimized VQ-AE and Application
Figure 3: Illustration of cascade pipeline: (a) displays the overview of Wav2Lip; (b) displays the latent denosing process of SOVQAE and the mechanism of regularization loss. Please zoom in for better visualization.
4.1 Space Optimized VQ-AE
@@
According to Theorem 3.3, perfect denoising is achievable when $\|\mathcal{N}\|_{F}<\frac{d_{C}-2\gamma}{2L_{\epsilon}}$. Hence our goal is increasing the NRoUB thus to enhance the noise robustness of VQ-AE. Note that $\gamma$ is much close to zero, because $\gamma$ is optimized via VQ loss (see Eq. 8) in training. Hence, it seems optimal to maximize $d_{C}$ while minimizing $L_{\epsilon}$. However, this is theoretically infeasible, because $L_{\epsilon}$ is positively correlated with the operation norm of convolutional kernels (according to our proof of Theorem 3.2), implying that minimizing $L_{\epsilon}$ equates to L2/L1 parameter regularization goodfellow2016deep . Such regularization would globally compress the latent space distribution, consequently compressing the distribution of anchors in the codebook $\mathbf{C}$ and reducing $d_{C}$, thus to degrade NRoUB, as demonstrated in Tab. 3.
@@
Therefore, the collect optimization is that increasing $d_{C}$ while maintaining $L_{\epsilon}$. In experiments, we find a practical solution is to locally extend the $d_{C}$ with a given lower bound. Thus, the codebook regularization loss we propose is:
@@
where $\theta$ represents the given lower bound. This regularization strategy effectively improves local noise robustness while maintaining global robustness, which will be validated via ablation studies. In addition to this regularization loss, the complete loss function of SOVQAE comprises L2 reconstruction loss, VQ lossesser2021taming , perceptual loss, and GAN loss with a patch-based discriminatorisola2017image . The VQ loss is following:
@@
where $x$ is the input image, $\hat{x}=\delta(g_{\textbf{C}}(\epsilon(x)))$ is the reconstructed image, $\text{sg}(\cdot)$ denotes the stop-gradient operation followingesser2021taming . The later two components of Eq. 8 are utilized to optimized the $\gamma$.
4.2 Cascade ID-specific THG pipeline
The Choice of Base Model. In ID-specific THG, an important demanding is the generality with Out-of-Domain (OOD) audios. While end-to-end methods ye2023geneface++ ; peng2023synctalk make significant progress on it, they still have gap with great OOD audio generality, as their limited training data. Furthermore, while Wav2Lip prajwal2020lip has been proposed for years, it still outperforms with both recent one-shot tan2024edtalk ; kim2025moditalker and ID-specific xie2025pointtalk ; yu2024gaussiantalker methods and run in real time. It is an ideal lip-sync prior base model for a post-processing framework. As shown in Tab. 1, this combination makes outstanding lip-sync performance on OOD audios, validating our choice of base model.
Cascade Pipeline. As shown in Fig. 3, Our SOVQAE is Cascaded with pretrained Wav2Lip, and we kindly refer readers to obtain more details from prajwal2020lip due to the limited space. To adapt with outputs of Wav2Lip, we employ face detection using S3FD zhang2017s3fd to extract facial regions from a talking head video of a specific identity. These extracted regions are subsequently resized to $256\times 256$ for training SOVQAE. This strategy allows the codebook to concentrate on capturing the high-frequency facial components. In inference, for given Low resolution faces ($96\times 96$) from Wav2Lip, we upscale them to $256\times 256$ before subjecting them to SQVQAE.
Efficiency and Plug-and-Play. With the half precision training and inference, our cascade pipeline achieves 30 FPS inference speed on consumer GPU, which demonstrates its feasibility in industry. Total training takes 5 GPU hours on NVIDIA 4090, which is also comparable with end-to-end counterparts. As shown in Fig. 10, trained SOVQAE can also adapt for different base models, since it works as a denoiser.
Figure 4: The comparative visualization presents a curated selection of discrete keyframes and corresponding sub-frames of lower-half faces. Please zoom in for better visualization.
@@ 5 Experiments
Dataset. For comparative analysis, we employ the same set of meticulously edited video sequences utilized in previous works guo2021ad ; li2023efficient ; ye2023geneface , encompassing both English and French dialogues. The average duration of these videos is approximately 8,843 frames. We refer to this collection as the Usual Dataset to differentiate it from our proprietary HF videos. These videos are characterized by their rich textures and high resolution ($1024\times 1024$). The average length of these videos is around 11,000 frames. More details of them can be found in Appendix. All videos are centered on individual characters and captured at a frame rate of 25 FPS. Consistent with prior research, we adopt a train-to-test split ratio of $10:1$. We empirically set $\theta=1$. Moreover, to assess the generalizability of our methods, we conduct additional tests using a set of out-of-domain (OOD) audio. This OOD audio comprises 10 diverse voice samples, spanning various languages and genders. We place more implementation details in Appendix, due to the limited space.
@@
Comparison Baselines. We conduct a comparative analysis of our method with a total of three one-shot approaches, which include Wav2Lip prajwal2020lip , DINet zhang2023dinet , IP-LAP zhong2023identity , and three ID-specific THG methods, namely SyncTalk peng2023synctalk , GeneFace ye2023geneface and GeneFace++ ye2023geneface++ . Furthermore, to underscore the significance of our identity-specific face restoration concept, we establish two post-processing baselines using Codeformer zhou2022towards and GFPGAN wang2021towards .
@@
Evaluation Metrics. To assess the quality of the generated images, we employ the Peak Signal-to-Noise Ratio (PSNR) and the Frechet Inception Distance (FID) heusel2017gans as metrics. For evaluating lip synchronization, we utilize the Lip Sync Error Confidence (LSE-C) and Lip Sync Error Distance (LSE-D) scores.
5.1 Qualitative Comparison
To provide a more intuitive assessment of image quality, we present a comparative analysis of our method alongside alternative approaches in Fig. 4. Specifically, we contrast our method with both end-to-end methodologies, including Wav2Lip prajwal2020lip , SyncTalk peng2023synctalk , IP-LAP zhong2023identity , and DINet zhang2023dinet , as well as two post-processing techniques, GFPGAN wang2021towards and CodeFormer zhou2022towards .
Figure 5: Comparison on HF videos. Except for the collapse and lip error (see red boxes), NeRF-based methods also has limitation in stability of face scale. See supplemental videos for more details.
In summary, the qualitative comparison underscores that our method surpasses all compared methods in terms of facial texture and lip-synchronization. In Fig. 4, our method outperforms SyncTalk, a state-of-the-art NeRF-based approach, by better capturing high-frequency facial details. SyncTalk’s struggle to maintain these details, particularly in dynamic scenes, results in blurred mouth and teeth regions. This is attributed to NeRF’s limited expressivity for high-frequency details. In contrast, our method, utilizing SOVQAE, effectively restores individual-specific textures, enhancing fidelity. The 7th and 8th columns of the figure, where our method’s identity-specific restoration approach is superior to the two post-processing baselines, which alter the subject’s identity due to information leakage from a large OOD faces dataset karras2019style . This underscores the necessity of our identity-specific restoration approach. Besides, in more challenging HF videos, our method has more stable performance. In contrast, NeRF-based methods occur geometry collapse (see the 3-th cloumn in Fig.5) and lip error, when scale their resolution to higher (from 512 to 1024).
@@ 5.2 Quantitative Comparison
For a more exhaustive evaluation, our quantitative assessment is divided into two segments: 1) Video quality and lip-synchronization comparison, which evaluates image quality frame-by-frame and lip-synchronization performance under original audio-driven conditions; 2) Out-of-Distribution (OOD) audio-driven comparison, where we contrast lip-synchronization with SOTA methods to underscore the superior cross-audio generality of our approach, a critical application scenario.
Comparison with End-to-End Methods. As detailed in Tab. 1, on the Usual Dataset, our method secures the highest scores for both PSNR (38.9213 ) and FID (3.5502), signifying that we achieve state-of-the-art image quality for generated videos. This advancement is credited to the SOVQAE, which retains high-frequency texture details in the codebook and consistently recovers facial textures from low-quality faces. In terms of lip-synchronization, we outperform all methods (excluding our base model), a testament to Wav2Lip’s robust audio-lip alignment capability. Although our method moderately reduces the lip-synchronization performance of the base model, the substantial enhancement in visual quality justifies this trade-off as valuable and insightful. Similar outcomes are observed in the HF videos.
Comparison with Post-Process Methods. As illustrated in Tab. 1, our method surpasses two post-process baselines in video quality. This advantage stems from the fact that Codeformer and GFPGAN are tailored for blind face restoration, and their face prior learning from out-of-distribution multi-face datasets results in information leakage, leading to identity errors. Moreover, their image-by-image training strategy is not optimal for video scenarios, causing frame-wise flickering. Although they maintain the lip-synchronization performance of the base model, their subpar visual fidelity limits their applicability to ID-specific THG. Furthermore, for fair comparison, we also finetune GFPGAN model in ID-specific videos. As shown in Fig. 6, while fine-tuning improves its performance in visual quality, its synchronization decreases. On the other hand, our method has least GPU need (5 GPU@4090 hours), best visual quality and comparable synchronization. These findings strongly affirm the efficacy and efficiency of our cascade pipeline for ID-specific THG.
Comparison on OOD Audio Driving. To evaluate our method’s voice generalization, we compared it with recent SOTA models on the OOD audio-driving task. As shown in Tab. 1, Our method achieves second best in both LSE-C and LSE-D metrics, scoring 6.5521 in LSE-C and 7.6231 in LSE-D, demonstrating strong generalization across diverse voices. It significantly outperforms other methods such as SyncTalk, GeneFace++, IP-LAP, and DINet, and is slightly behind Wav2Lip.
Table 1: The quantitative experiments of different methods on Usual Dataset and HF videos, and the OOD audio-driven experiment. * indicates post-process baselines combined with Wav2Lip.
Method Usual Dataset HF videos OOD-audio
@@
Wav2Lipprajwal2020lip 34.0979 10.8503 9.077 5.8769 34.6593 17.5281 8.8496 6.8712 7.1258 7.3521
DINetzhang2023dinet 33.9108 9.8385 7.1951 7.4343 34.4849 10.5749 8.3541 7.2340 5.2310 8.5922
IP-LAPzhong2023identity 34.7263 10.1703 5.1736 8.8160 35.1595 9.9438 8.2604 7.1838 4.3759 9.0596
GeneFaceye2023geneface 24.8165 21.7084 5.195 - - - - - - -
GeneFace++ye2023geneface++ 31.1164 20.5506 6.8916 7.5014 33.9692 22.3928 4.6112 11.0401 5.3445 8.3927
SyncTalkpeng2023synctalk 36.0574 6.4855 7.054 7.282 39.4050 6.5370 7.816 7.715 5.1063 8.0474
CodeFormer*zhou2022towards 33.2441 26.3567 9.1739 5.7720 33.7958 16.7215 7.9409 7.2582 - -
GFPGAN*wang2021towards 33.9803 16.7267 9.0417 5.8144 33.7780 17.2715 8.7940 6.5193 - -
Ours 38.9213 3.5502 8.1136 6.3402 39.6254 4.3525 8.2392 6.8238 6.5521 7.6231
@@
Table 2: Plug-and-play experiments in Usual Dataset.
@@
GeneFace++ 31.11 20.55 6.89 7.50
GeneFace++ + SOVQAE 33.27 18.21 6.42 8.22
SyncTalk (CVPR 2024) 36.06 6.49 7.05 7.68
SyncTalk + SOVQAE 37.16 5.12 6.58 7.88
Ours 38.92 3.55 8.11 6.34
Plug-and-Play Performance. To demonstrate the plug-and-play performance of SOVQAE, we integrated SOVQAE into different ID-specifc base models. As shown in Tab.2, SOVQAE improves visual quality metrics across different ID-specific base models, confirming its versatility and effectiveness. Notably, we achieve the best overall performance by leveraging the exceptional audio-lip synchronization capabilities of Wav2Lip. This superior synchronization positively impacts PSNR and FID values, thereby enhancing overall video quality and ensuring ours top performance.
These observations suggest that our method effectively inherits the SOTA audio generalization capability of Wav2Lip and validate the feasibility of our cascade pipeline.